Find the equation of the line that passes through the points (-2,7 and 6,-9).
Answer by Alan3354(69443)

|+x +y +1|
|-2 +7 +1| = 0
|+6 -9 +1|
------
x*(7+9) - y*(-2-6) + 1*(18-42) = 0
16x + 8y = 24
2x + y = 3